The area of a rectangular playground is 16 square yards. The length of the playground is 4 times longer than its width. Find the length and width of the playground
Answer by Euler(11)

Hi!
Okay, suppose the width here is x, so the length would be 4x.
x*4x=16
4x^2=16
x^2=4
x=2
Width=2 Length=8
Recheck
2*4(2)=16
2*8=16
16=16 :)